OpenStudy (anonymous):

Need some help here, and I need to show my work and really have no clue as to how i do these any help is appreciated. :) Solve each equation by factoring. 1. x^2 + 14x – 72 = 0 2. n^2 + 4n = 0 3. 2x2 + 11x + 12 = 0 4. g^2 + 10g + 25 = 0 Solve each equation by using the Quadratic Formula. 5. 25b^2 – 20b = 0 6. 3s^2 + 31x + 36 = 0 7. k + 2 = -3k^2 8. 2v^2 + 7x = 15 9. 2p^2 + 3p – 5 = 0 10. x^2 – 5x = -4 Any help??

OpenStudy (anonymous):

in #1 is that x sqaured

OpenStudy (anonymous):

yes, my bad meant to put a carrot in there.

OpenStudy (anonymous):

its good

OpenStudy (anonymous):

(x+18)(x-4) is #1

OpenStudy (anonymous):

do u mind if i phone a friend :)

OpenStudy (anonymous):

Don't mind a bit :)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@david111

OpenStudy (anonymous):

1.x=-18,4

OpenStudy (anonymous):

2.n=0,−4

OpenStudy (anonymous):

3.x=−4,−3/2

OpenStudy (anonymous):

4.g=−5

OpenStudy (anonymous):

well ill leave you two alone then

OpenStudy (anonymous):

5.b=4/5,0

OpenStudy (anonymous):

6.x=−3s^2/31−36/31

OpenStudy (anonymous):

7.1/6(−1+i23√)

OpenStudy (anonymous):

8.v=√15/2−7x/2

OpenStudy (anonymous):

9.p=−5/2

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

10.x=4,1

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

i told u she is good

OpenStudy (anonymous):

np